# A mean-field theory for heterogeneous random growth with redistribution

### v2

@@ Abstract
For $\sigma=0$, Eq. (2) can be rewritten in a matrix form $\dot{\bf x}=\mathbb{M}{\bf x}$, with $\mathbb{M}_{ij}=(m_{i}-\varphi)\delta_{ij}+{\varphi}/{N}$. Noting that $\mathbb{M}$ is a diagonal matrix plus a rank one perturbation, the characteristic polynomial can easily be computed (see Appendix) and leads, for $\varphi>0$, to the following equation for the eigenvalues $\gamma_{\alpha}$
@@
We now study (3). For fast redistribution, $\varphi\to+\infty$, the overall growth rate converges to the average of the local growth rate over all $i$, i.e one finds (see Appendix)
@@
The first term is expected since the population keeps hopping between all cities, and idiosyncratic effects are averaged out. The correction is positive: decreasing redistribution always increases the overall growth rate. Indeed we show from (3), see Appendix, that $1\leq\partial_{\varphi}\gamma<0$.
@@
The above results generalize to a larger class of distributions with finite support, and with $\rho(m)\sim(m_{>}-m)^{\psi-1}$ close to the upper edge, with $\psi=3/2$ (resp. $1/2$) for the Wigner (resp. arc-sine) case. Whenever $\psi>1$, there is a localisation transition at $\varphi_{c}=1/G(m_{>})$. For $\varphi>\varphi_{c}$, the growth rate $\gamma$ is the unique solution of Eq. (5), whereas for $\varphi<\varphi_{c}$, the population localises and the growth rate is given again by $\gamma=m_{>}-\varphi$. In the case $\psi\leq 1$ on the other hand, $G(m_{>})=\infty$ and the population remains delocalised as soon as $\varphi>0$. The resulting $\gamma(\varphi)$ is checked numerically in Fig. 2 in the Appendix.
Note that, interestingly, an equation similar to Eq. (5) appeared very recently Gueneau2025 in a model of a 1D Brownian particle that resets its diffusion coefficient $D$ with rate $r=\varphi$, to a random value distributed according to $\rho(D)$. The analog of our transition for $N\to+\infty$ first, i.e. for an infinite pool of values $D_{i}$, was observed in the large time limit by the authors of Gueneau2025 . For a finite pool $\{D_{i}\}_{i\in[1,N]}$, we thus predict the same condensation transition (dominance of $D_{\rm max}$) as obtained here, see Appendix.
@@
meaning that indeed, when ${\varphi}<{\varphi_{c}}$ a finite fraction of the population is in a single “state” occupation , as in the Bose condensation problem. This fraction vanishes continuously as ${\varphi}\to{\varphi_{c}}^{-}$, and remains zero in the delocalised phase, see Fig. 3 in the Appendix.
@@
We now turn to the case where $\rho(m)$ has an infinite support, with $m_{>}=+\infty$. In order to be specific, we consider the following family of generalized Gaussians: $\rho(m)=\exp{(-(m^{2}/2\Sigma^{2})^{b})}/Z_{b}$ where $b>0$ and $Z_{b}$ an appropriate normalisation. The Gaussian case corresponds to $b=1$, whereas $b=1/2$ is the double Laplace distribution. From extreme value theorems, we know that the largest value of $m$ after $N\gg 1$ independent draws is $m_{1}\sim\Sigma\sqrt{2(\log N)^{1/b}}$. To keep the maximum growth rate finite as $N$ goes to infinity, we choose $\Sigma=\Sigma_{0}/\sqrt{2(\log N)^{1/b}}$, such that $m_{1}\approx\Sigma_{0}$.
@@
with the self-consistent condition that $Z(t)\approx N\Gamma e^{\gamma t}$ at large times, coming from the fact that $\sum_{i}x_{i}(t)=N\overline{x}(t)$.
@@
$\mu=\frac{\sigma^{4}-\Sigma_{0}^{2}}{\sigma^{2}\sqrt{\sigma^{4}+\Sigma_{0}^{2}}}.$ (17)
Using the mapping to the REM, we predict that the distribution of $p_{\max}=\max_{i}x_{i}(t)/\bar{x}(t)$ decays as $\sim 1/p_{\max}^{1+\mu}$, as checked in Fig.5. For $\sigma<\Sigma_{0}$, $I_{2}$ starts diverging when $\varphi+\gamma+\sigma^{2}/2\to\Sigma_{0}^{+}$, giving a contribution $\sim[N(\varphi+\sigma^{2}/2+\gamma-\Sigma_{0})]^{-1}$ which becomes of order one. This is the analogue of the calculation of $p_{1}$ above, which becomes of order unity in the localised phase. Hence in that phase, $\gamma=\gamma_{\rm loc}=\varphi_{c}^{I-II}-\varphi$ with $\varphi_{c}^{I-II}:=\Sigma_{0}-{\sigma^{2}}/{2}$.
@@
Hence, our REM analysis predicts a rather rich phase diagram with three phases, summarized in Fig. 1 in the plane $(\varphi/\Sigma_{0},\sigma^{2}/\Sigma_{0})$. When $\sigma^{2}=0$, we recover the results above, Eq. (9), using a different framework, which emphasizes the deep connections of our problem with the REM. But as $\sigma^{2}$ becomes large enough, the localised phase observed for $\varphi<\Sigma_{0}$ disappears and becomes delocalised with $\gamma=0$. When $0<\varphi<\Sigma_{0}/2$, the system first goes through a partially localised phase with $\gamma=\gamma_{1}$ before becoming fully delocalised. In particular, our result recovers the fact that there is no localisation for large $N$ when $\Sigma_{0}=0$ BouchaudMezardCondensation2000 . [A more refined analysis shows however that there is a crossover between a localised region and delocalised region when $\log N\sim\sigma^{2}/2\varphi$ footnotecrossover , which thus becomes relevant when $\varphi\to 0$.] We have performed numerical checks of these results. We have clear evidence of a localised phase with the predicted value for $\gamma_{\rm loc}$, see Fig. 4 in the Appendix. We also confirm the presence of the partially localised phase, although the precise determination of $\gamma_{1}$ is made difficult by $1/\log{N}$ corrections. Note that our model corresponds to the mean-field limit of the model considered in Hur , which did not mention the existence nor the nature of the three phases reported here. Many directions are left to explore, such as the influence of heterogeneities in the variance $\sigma^{2}$ across individuals or of the network topology on the phase diagram of Fig. 1.
@@
In conclusion, heterogeneous mean growth rates lead to much stronger level of inequalities than volatile but homogeneous growth rates. In the latter case, uniform redistribution, however small, always leads to moderate inequalities, characterised by fat tails but with a finite mean BouchaudMezardCondensation2000 , as we mentioned in the introduction. In contrast, when mean growth rates are heterogeneous, a minimum level of redistribution $\varphi_{\text{c}}$ is needed to ensure that extreme concentration does not occur. However, this comes at a cost since the corresponding global growth rate $\gamma$ is substantially reduced from its maximum value, dominated by outsiders. When both effects are present, a sufficient level of temporal fluctuations in growth rates can mitigate extreme concentrations (see Fig. 1). In the context of wealth inequalities, this means that persistent differences of “skill” or ”rents” (i.e. different $m_{i}$s) can only lead to extreme concentrations (or “oligarchies” as coined in Boghosian2019 ), except if a sufficient wealth tax rate $\varphi>\varphi_{c}$ is introduced as a redistribution tool. This is still the case when success is the result of a mix of “skill” and “luck” Pluchino ; Hur , albeit with a lower value of $\varphi_{c}$. When the luck component $\sigma^{2}$ is large enough (i.e. $\Sigma_{0}$), a smaller wealth tax rate $\varphi\geq\Sigma_{0}^{2}/2\sigma^{2}$ is enough to prevent extreme concentration. In our highly stylized setting, accidental “bad streaks” are enough to avoid a fraction of total wealth being concentrated in the hands of a few individuals. Even in the partially localised phase, interestingly, the “happy few” are not fixed in time: all individuals get lucky one day or another.

### v3

@@ Abstract
For $\sigma=0$, Eq. (2) can be rewritten in a matrix form $\dot{\bf x}=\mathbb{M}{\bf x}$, with $\mathbb{M}_{ij}=(m_{i}-\varphi)\delta_{ij}+{\varphi}/{N}$. Noting that $\mathbb{M}$ is a diagonal matrix plus a rank one perturbation, the characteristic polynomial can easily be computed (see SM SM ) and leads, for $\varphi>0$, to the following equation for the eigenvalues $\gamma_{\alpha}$
@@
We now study (3). For fast redistribution, $\varphi\to+\infty$, the overall growth rate converges to the average of the local growth rate over all $i$, i.e one finds (see SM SM )
@@
The first term is expected since the population keeps hopping between all cities, and idiosyncratic effects are averaged out. The correction is positive: decreasing redistribution always increases the overall growth rate. Indeed we show from (3), see SM SM , that $1\leq\partial_{\varphi}\gamma<0$.
@@
The above results generalize to a larger class of distributions with finite support, and with $\rho(m)\sim(m_{>}-m)^{\psi-1}$ close to the upper edge, with $\psi=3/2$ (resp. $1/2$) for the Wigner (resp. arc-sine) case. Whenever $\psi>1$, there is a localisation transition at $\varphi_{c}=1/G(m_{>})$. For $\varphi>\varphi_{c}$, the growth rate $\gamma$ is the unique solution of Eq. (5), whereas for $\varphi<\varphi_{c}$, the population localises and the growth rate is given again by $\gamma=m_{>}-\varphi$. In the case $\psi\leq 1$ on the other hand, $G(m_{>})=\infty$ and the population remains delocalised as soon as $\varphi>0$. The resulting $\gamma(\varphi)$ is checked numerically in Fig.S3 of the SM SM .
Note that, interestingly, an equation similar to Eq. (5) appeared very recently Gueneau2025 in a model of a 1D Brownian particle that resets its diffusion coefficient $D$ with rate $r=\varphi$, to a random value distributed according to $\rho(D)$. The analog of our transition for $N\to+\infty$ first, i.e. for an infinite pool of values $D_{i}$, was observed in the large time limit by the authors of Gueneau2025 . For a finite pool $\{D_{i}\}_{i\in[1,N]}$, we thus predict the same condensation transition (dominance of $D_{\rm max}$) as obtained here, see Section C in SM SM .
@@
meaning that indeed, when ${\varphi}<{\varphi_{c}}$ a finite fraction of the population is in a single “state” occupation , as in the Bose condensation problem. This fraction vanishes continuously as ${\varphi}\to{\varphi_{c}}^{-}$, and remains zero in the delocalised phase, see Fig. S3 in the SM SM .
@@
We now turn to the case where $\rho(m)$ has an infinite support, with $m_{>}=+\infty$. In order to be specific, we consider the following family of generalized Gaussians: $\rho(m)=\exp{(-(m^{2}/2\Sigma^{2})^{b})}/Z_{b}$ where $b>0$ and $Z_{b}$ an appropriate normalisation. The Gaussian case corresponds to $b=1$, whereas $b=1/2$ is the double Laplace distribution. From extreme value theorems, we know that the largest value of $m$ after $N\gg 1$ independent draws is $m_{1}=m_{1}(N)\sim\Sigma\sqrt{2(\log N)^{1/b}}$. To keep the maximum growth rate finite as $N$ goes to infinity, we choose $\Sigma=\Sigma_{0}/\sqrt{2(\log N)^{1/b}}$, such that $m_{1}\approx\Sigma_{0}$.
@@
with the self-consistent condition that $Z(t)\approx N\Gamma e^{\gamma t}$ at large times, coming from the fact that $\sum_{i}x_{i}(t)=N\overline{x}(t)$. footnotefluct
@@
$\mu=1-\frac{\Sigma_{0}^{2}}{\sigma^{4}}$ (17)
Using the mapping to the REM, we predict that the distribution of $p_{\max}=\max_{i}x_{i}(t)/\bar{x}(t)$ decays as $\sim 1/p_{\max}^{1+\mu}$, as checked in Fig.S3 of the SM SM . For $\sigma<\Sigma_{0}$, $I_{2}$ starts diverging when $\varphi+\gamma+\sigma^{2}/2\to\Sigma_{0}^{+}$, giving a contribution $\sim[N(\varphi+\sigma^{2}/2+\gamma-\Sigma_{0})]^{-1}$ which becomes of order one. This is the analogue of the calculation of $p_{1}$ above, which becomes of order unity in the localised phase. Hence in that phase, $\gamma=\gamma_{\rm loc}=\varphi_{c}^{I-II}-\varphi$ with $\varphi_{c}^{I-II}:=\Sigma_{0}-{\sigma^{2}}/{2}$. We test this prediction in the inset of Fig. 2 while the main plot shows the distribution of $p_{\rm max}=\max_{i}p_{i}$. As $N$ becomes large, $p_{\rm max}$ is $O(1)$ with fluctuations of the same order. Note that $\gamma_{\rm loc}=\Sigma_{0}-\varphi-\frac{\sigma^{2}}{2}=\lim_{N\to+\infty}m_{1}(N)-\varphi-\frac{\sigma^{2}}{2}$ corresponds to the asymptotic growth rate $m_{1}(N)$ of the site where a finite fraction of the population concentrates. This allows us to predict finite size corrections.
@@
Figure 2: Main: Plot of the probability distribution of $p_{\rm max}=\max_{i}x_{i}/\sum_{j}x_{j}$ for various NN for $\varphi=0.2$ and $\sigma=0.3$. It shows that the system is localised for these values of the parameters. Inset: corresponding effective growth rates measured as $\langle\log{\frac{\bar{x}(t)}{\bar{x}(t_{0})}}\rangle/(t-t_{0})$ for $t_{0}=40$ (continuous line), compared to the prediction $\gamma=m_{1}-\varphi-\frac{\sigma^{2}}{2}$ (dashed lines). The value of $m_{1}=m_{1}(N)$ converges to $\Sigma_{0}=1$ with logarithmic corrections.footnotefigpmax
@@
Hence, our REM analysis predicts a rather rich phase diagram with three phases, summarized in Fig. 1 in the plane $(\varphi/\Sigma_{0},\sigma^{2}/\Sigma_{0})$. When $\sigma^{2}=0$, we recover the results above, Eq. (9), using a different framework, which emphasizes the deep connections of our problem with the REM. But as $\sigma^{2}$ becomes large enough, the localised phase observed for $\varphi<\Sigma_{0}$ disappears and becomes delocalised with $\gamma=0$. When $0<\varphi<\Sigma_{0}/2$, the system first goes through a partially localised phase with $\gamma=\gamma_{1}$ before becoming fully delocalised. In particular, our result recovers the fact that there is no localisation for large $N$ when $\Sigma_{0}=0$ BouchaudMezardCondensation2000 . [A more refined analysis shows however that there is a crossover between a localised region and delocalised region when $\log N\sim\sigma^{2}/2\varphi$ footnotecrossover , which thus becomes relevant when $\varphi\to 0$.] We have performed numerical checks of these results. We have clear evidence of a localised phase with the predicted value for $\gamma_{\rm loc}$, see Fig. 2 in the SM SM . We also confirm the presence of the partially localised phase, although the precise determination of $\gamma_{1}$ is made difficult by $1/\log{N}$ corrections. Note that our model corresponds to the mean-field limit of the model considered in Hur , which did not mention the existence nor the nature of the three phases reported here. Many directions are left to explore, such as the influence of heterogeneities in the variance $\sigma^{2}$ across individuals or of the network topology on the phase diagram of Fig. 1.
@@
In conclusion, heterogeneous mean growth rates lead to much stronger level of inequalities than volatile but homogeneous growth rates. In the latter case, uniform redistribution, however small, always leads to moderate inequalities, characterised by fat tails but with a finite mean BouchaudMezardCondensation2000 , as we mentioned in the introduction. In contrast, when mean growth rates are heterogeneous, a minimum level of redistribution $\varphi_{\text{c}}$ is needed to ensure that extreme concentration does not occur. However, this comes at a cost since the corresponding global growth rate $\gamma$ is substantially reduced from its maximum value, dominated by outliers. When both effects are present, a sufficient level of temporal fluctuations in growth rates can mitigate extreme concentrations (see Fig. 1). In the context of wealth inequalities, this means that persistent differences of “skill” or ”rents” (i.e. different $m_{i}$s) can only lead to extreme concentrations (or “oligarchies” as coined in Boghosian2019 ), except if a sufficient wealth tax rate $\varphi>\varphi_{c}$ is introduced as a redistribution tool. This is still the case when success is the result of a mix of “skill” and “luck” Pluchino ; Hur , albeit with a lower value of $\varphi_{c}$. When the luck component $\sigma^{2}$ is large enough (i.e. $\Sigma_{0}$), a smaller wealth tax rate $\varphi\geq\Sigma_{0}^{2}/2\sigma^{2}$ is enough to prevent extreme concentration. In our highly stylized setting, accidental “bad streaks” are enough to avoid a fraction of total wealth being concentrated in the hands of a few individuals. Even in the partially localised phase, interestingly, the “happy few” are not fixed in time: individuals may get lucky one day or another. footnoteconcl